OpenStudy (anonymous):

I don't really get this. Question: Use the rules for exponents and roots to prove that 512^2/3 = 16^3/2

OpenStudy (anonymous):

Do you mean a rigorous formal proof?

OpenStudy (anonymous):

I don't really know. This is what the question was and I don't get it.

OpenStudy (anonymous):

Do you understand how to use logarithms?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Ah, you will want to understand natural logarithms for this question, I can explain it.

OpenStudy (anonymous):

\[y = \log_b x\] \[x = b^y\]

OpenStudy (anonymous):

If we take the natural logarithm of \[512^{\frac{2}{3}}\] \[\frac{2}{3}\ln(512)\]

OpenStudy (anonymous):

We can pull the exponent out

OpenStudy (anonymous):

So, is that a cube root? That we ended up with?

OpenStudy (precal):